OpenStudy (anonymous):

By how much does a^3 - 2a exceed a^2 + a - 6?

OpenStudy (anonymous):

since a^3 - 2a exceed the other no so it is bigger... = (a^3 - 2a) - (a^2 + a -6) = a^3 - 2a - a^2 - a + 6 = a^3 - a^2 - 3a + 6

OpenStudy (saifoo.khan):

Yay, Myinin